# ISRE-0001-2024.00

### v1

@@ 3.3 Theoretical Analysis
In this section, we examine the sufficient conditions under which $\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g}$ is a consistent estimator for $\bm{D}_{\tilde{y}\neq y^{*}|g}$. We show that, 1) when there is no model noise, and the model $f_{i}:\bm{x}_{g_{k}}\rightarrow\hat{p}(\bm{x}_{g_{k}})$ produces ideal predicted probabilities $\hat{p}(\bm{x}_{g_{k}})$ (defined in Condition 1), or 2) when the model is imperfect and produces per-instance diffracted noisy predicted probabilities (defined in Condition 2), DeCoLe exactly identifies mislabeled instances in a dataset with label bias. To disambiguate label noise and model noise, we consider two levels of assumption on model noise and the predicted probability $\hat{p}(\bm{x}_{g_{k}})$, and we show that under both ideal and noisy $\hat{p}(\bm{x}_{g_{k}})$, DeCoLe captures erroneous instances in a dataset with label bias and consistently estimates the set of mislabeled instances $\{\bm{D}_{\tilde{y}\neq y^{*}|g}\}$. We demonstrate this in the series of theorems below. We also note that we complement these theoretical guarantees with empirical results in Section 4, where we show that the proposed method yields state-of-the-art results even when the theoretical conditions identified in this section are violated.
@@
We begin by considering an ideal scenario that builds on a well-established assumption in the confident learning literature, where predicted probabilities are modeled to capture the class-conditional noise structure for detecting label noise (Northcutt et al. 2021). We extend this framework to account for group- and class-conditional noise, acknowledging that label noise can arise not only from class-dependent errors but also from systematic biases across groups. Building on the confident learning literature and under our extended framework, Condition 1 assumes that the predicted probabilities $\hat{p}(\bm{x}_{g_{k}})$ accurately reflect the underlying group- and class-conditional noise in the labels. We show that, under Condition 1, DeCoLe exactly detects mislabeled instances in a dataset with group and class conditional noise (label bias) and that $\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g}$ is a consistent estimator for $\bm{D}_{\tilde{y}\neq y^{*}|g}$.
@@
Let $\hat{p}(\tilde{y}=1;\bm{x}_{g_{k}},f_{k})$ denote the probability predicted by classifier $f_{k}$ that an instance $\bm{x}_{g_{k}}$ belongs to a positive class. And let $\hat{p}(\bm{x}_{g_{k}})$ be the shorthand writing for $\hat{p}(\tilde{y}=1;\bm{x}_{g_{k}},f_{k})$.
Condition 1 (Ideal Predicted Probability)
@@
When the predicted probability $\hat{p}(\bm{x}_{g_{k}})$ is ideal, we have $\hat{p}(\bm{x}_{g_{k}})=\hat{p}(\tilde{y}=1;\bm{x}_{g_{k}},f_{k})=\hat{p}(\tilde{y}=1;\bm{x}_{g_{k}}\in\bm{D}_{y^{*}=r,g=g_{k}},f_{k})=p(\tilde{y}=1;\bm{x}_{g_{k}}\in\bm{D}_{y^{*}=r,g=g_{k}},f_{k})=p(\tilde{y}=1|y^{*}=r,g=g_{k})=p(\bm{x}_{g_{k}})$. Essentially, ideal predicted probability $\hat{p}(\bm{x}_{g_{k}})$ match the noise rates $p(\tilde{y}=1|y^{*}=r,g=g_{k})$ corresponding to true label $r$ for $\bm{x}_{g_{k}}$.
Theorem 1
@@
For a dataset $\bm{D}\coloneqq(\bm{x},\tilde{y})^{n}$ with group- and class-conditional label noise $\pi_{0,{g_{k}}}=p(\tilde{y}=0|y^{*}=1,g=g_{k})<0.5$ and $\pi_{1,{g_{k}}}=p(\tilde{y}=1|y^{*}=0,g=g_{k})<0.5$, if the classifiers $f_{k}$ produce ideal predicted probabilities, then the set of detected mislabeled instances $\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g}$ is a consistent estimator for $\bm{D}_{\tilde{y}\neq y^{*}|g}$.
@@
While Condition 1 provides a useful sanity check, we now relax this assumption to capture a more nuanced and realistic scenario where the noisy predicted probability $\hat{p}(\bm{x}_{g_{k}})$ can vary at the instance level. This assumption is well-established and widely adopted in prior studies (e.g. Jiang et al. (2020), Northcutt et al. (2021)). We extend this perspective to explicitly incorporate both group- and class-conditional noise structures. Under this extended framework, we show that DeCoLe successfully identifies mislabeled instances in a dataset with label bias and that $\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g}$ remains a consistent estimator for $\bm{D}_{\tilde{y}\neq y^{*}|g}$.
Condition 2 (Per-instance diffracted (Noisy) Predicted Probability)
@@
The predicted probability $\hat{p}(\bm{x}_{g_{k}})$ provided by model $f_{k}$ is per-instance diffracted if it follows the relationship $\hat{p}(\bm{x}_{g_{k}})=p(\bm{x}_{g_{k}})+\epsilon_{\bm{x}_{g_{k}}}$ where the noise term $\epsilon_{\bm{x}_{g_{k}}}$ is drawn from the following distribution: $\epsilon_{\bm{x}_{g_{k}}}\sim\mathcal{U}[\epsilon_{k}+\text{LB}^{*}_{g_{k}}-p(\bm{x}_{g_{k}}),\epsilon_{k}-\text{LB}^{*}_{g_{k}}+p(\bm{x}_{g_{k}})]\hskip 1.99997pt$ when $p(\bm{x}_{g_{k}})>\frac{1}{2}$; and $\epsilon_{\bm{x}_{g_{k}}}\sim\mathcal{U}[\epsilon_{k}-\text{UB}^{*}_{g_{k}}+p(\bm{x}_{g_{k}}),\epsilon_{k}+\text{UB}^{*}_{g_{k}}-p(\bm{x}_{g_{k}})]$ when $p(\bm{x}_{g_{k}})<\frac{1}{2}$. Here, $\mathcal{U}$ denotes a uniform distribution, and $\epsilon_{k}=\mathbb{E}_{\bm{x_{g_{k}}}}[\epsilon_{\bm{x}_{g_{k}}}]$, where $\epsilon_{\bm{x}_{g_{k}}}$ represents the deviation from the ideal predicted probability. $\text{LB}^{*}_{g_{k}}$ and $\text{UB}^{*}_{g_{k}}$ denote the value of $\text{LB}_{g_{k}}$ and $\text{UB}_{g_{k}}$ under condition 1.
Theorem 2
@@
For a dataset $\bm{D}\coloneqq(\bm{x},\tilde{y})^{n}$, with group- and class-conditional label noise $\pi_{0,{g_{k}}}=p(\tilde{y}=0|y^{*}=1,g=g_{k})<0.5$ and $\pi_{1,{g_{k}}}=p(\tilde{y}=1|y^{*}=0,g=g_{k})<0.5$, if the classifiers $f_{k}:\bm{x}_{g_{k}}\rightarrow\hat{p}(\bm{x}_{g_{k}})$ yields per-instance diffracted predicted probabilities, then the set of detected mislabeled instances $\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g}$ is a consistent estimator for $\bm{D}_{\tilde{y}\neq y^{*}|g}$.
@@
The theoretical results establish that DeCoLe reliably performs bias-aware mislabeling detection under well-defined conditions. First, when the model produces ideal predicted probabilities, DeCoLe perfectly captures all erroneous labels. Then, relaxing this assumption, we consider the more realistic case where the predicted probabilities are per-instance diffracted (noisy), demonstrating that DeCoLe consistently identifies mislabeled instances. The detailed proofs of these theorems are provided in Appendix 7.
@@ 4.1 Evaluation Metrics
As illustrated in Figure 1, DeCoLe partitions the dataset into two disjoint subsets: (1) estimated mislabeled data, denoted by $\bm{\hat{D}}{\tilde{y}\neq y^{*}|g}$ and (2) estimated correctly labeled data, denoted by $\bm{\hat{D}}{\tilde{y}=y^{*}|g}$. This decomposition forms the foundation for evaluating DeCoLe’s performance. To quantify performance, we define four evaluation metrics:
@@
First, for data estimated to be mislabeled, we assess the proportion of truly mislabeled instances (denoted by $\bm{D}_{\tilde{y}\neq y^{*}|g}$) that are successfully identified for scrutiny. This corresponds to the overall recall of mislabeled instances, and is defined in Equation 5.
$\textit{Recall}_{\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g}}=\frac{|\bm{D}_{\tilde{y}\neq y^{*}|g}\cap\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g}|}{|\bm{D}_{\tilde{y}\neq y^{*}|g}|}$ (5)
@@
$\textit{Precision}_{\bm{\hat{D}}_{\tilde{y}=y^{*}|g}}=\frac{|\bm{D}_{\tilde{y}=y^{*}|g}\cap\bm{\hat{D}}_{\tilde{y}=y^{*}|g}|}{|\bm{\hat{D}}_{\tilde{y}=y^{*}|g}|}$ (6)
@@
$\textit{Recall}_{\bm{\hat{D}}_{\tilde{y}=c,y^{*}=\overline{c}|g}}=\frac{|\bm{D}_{\tilde{y}=c,y^{*}=\overline{c}|g}\cap\bm{\hat{D}}_{\tilde{y}=c,y^{*}=\overline{c}|g}|}{|\bm{D}_{\tilde{y}=c,y^{*}=\overline{c}|g}|}$ (7)
@@
$\textit{Precision}_{\bm{\hat{D}}_{\tilde{y}=y^{*}=c|g}}=\frac{|\bm{D}_{\tilde{y}=y^{*}=c|g}\cap\bm{\hat{D}}_{\tilde{y}=y^{*}=c|g}|}{|\bm{D}_{\tilde{y}=y^{*}=c|g}|}$ (8)
@@ 4.2.1 Data Generation
We create a synthetic dataset with group and class conditional noise rates, enabling full control over the relationship between observed labels $\tilde{y}$ and latent gold standard labels $y^{*}$. The synthetic population consists of $N=10{,}000$ instances, each associated with covariates $\bm{X}\in\mathbb{R}^{2}$, a binary group membership $g\in\{0,1\}$, a latent good standard label $y^{*}\in\{0,1\}$, and an observed label $\tilde{y}\in\{0,1\}$. We consider group imbalance, a frequent occurrence in real-world data (Mitchell et al. 2018), by assigning 70% of the population to a predominant group ($g=1$). Additionally, we account for differential subgroup validity (Hunter et al. 1979, De-Arteaga et al. 2022), a common phenomenon in which the relationship between covariates and target labels varies across groups. To do this, we draw covariates for different group and class combinations from bi-dimensional normal distributions with different means. We use the same standard deviation for all normal distributions. Further details about the sampling of $\bm{X}\in\mathbb{R}^{2}$ and $y^{*}\in\{0,1\}$ are provided in Appendix 6.
@@
To elucidate the phenomenon of differential subgroup validity and the rationale behind the decoupling component in DeCoLe, we visualize the linear classifiers that would be learned if the data is pooled together vs. if separate linear classifiers are learned for each group. It is easy to see that the classifiers for group $g_{0}$ only (Figure 3 (b)) and for group $g_{1}$ only (Figure 3 (c)) exhibit fundamental dissimilarity, being nearly orthogonal. Therefore, when we fit one classifier for both groups, as depicted in Figure 3 (a), it demonstrates the risks of differential subgroup validity. This phenomenon is closely related to Simpson’s Paradox (Simpson 1951). Note that the classification accuracy for the linear separator in Figure 3 (a) is notably higher for the majority group compared to the minority group. Furthermore, the classifier in Figure 3 (a) tends to misclassify positive instances in the minority group $g_{0}$ as negative and misclassify negative instances in the majority group $g_{1}$ as positive, reflecting how the classifier may learn and amplify bias in the data labels. If such a classifier is used as a building block of a mislabeling detection algorithm, then such bias will be encoded in it. As we will show in the following subsections, this results in poor mislabeling detection performance when the group-specific nature of predictive relationships and mislabeling patterns is not accounted for. The results also show how this issue is succesfully address by DeCoLe.

### v2

@@ 3.3   Theoretical Analysis
In this section, we examine the sufficient conditions under which $\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g}$ is a consistent estimator for $\boldsymbol{D}_{\tilde{y}\neq y^{*}|g}$. We show that, 1) when there is no model noise, and the model $f_{i}:\boldsymbol{x}_{g_{k}}\rightarrow\hat{p}(\boldsymbol{x}_{g_{k}})$ produces ideal predicted probabilities $\hat{p}(\boldsymbol{x}_{g_{k}})$ (defined in Condition 1), or 2) when the model is imperfect and produces per-instance diffracted noisy predicted probabilities (defined in Condition 2), DeCoLe exactly identifies mislabeled instances in a dataset with label bias. To disambiguate label noise and model noise, we consider two levels of assumption on model noise and the predicted probability $\hat{p}(\boldsymbol{x}_{g_{k}})$, and we show that under both ideal and noisy $\hat{p}(\boldsymbol{x}_{g_{k}})$, DeCoLe captures erroneous instances in a dataset with label bias and consistently estimates the set of mislabeled instances $\{\boldsymbol{D}_{\tilde{y}\neq y^{*}|g}\}$. We demonstrate this in the series of theorems below. We also note that we complement these theoretical guarantees with empirical results in Section 4, where we show that the proposed method yields state-of-the-art results even when the theoretical conditions identified in this section are violated.
@@
We begin by considering an ideal scenario that builds on a well-established assumption in the confident learning literature, where predicted probabilities are modeled to capture the class-conditional noise structure for detecting label noise (northcutt2021confident, ). We extend this framework to account for group- and class-conditional noise, acknowledging that label noise can arise not only from class-dependent errors but also from systematic biases across groups. Building on the confident learning literature and under our extended framework, Condition 1 assumes that the predicted probabilities $\hat{p}(\boldsymbol{x}_{g_{k}})$ accurately reflect the underlying group- and class-conditional noise in the labels. We show that, under Condition 1, DeCoLe exactly detects mislabeled instances in a dataset with group and class conditional noise (label bias) and that $\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g}$ is a consistent estimator for $\boldsymbol{D}_{\tilde{y}\neq y^{*}|g}$.
@@
Let $\hat{p}(\tilde{y}=1;\boldsymbol{x}_{g_{k}},f_{k})$ denote the probability predicted by classifier $f_{k}$ that an instance $\boldsymbol{x}_{g_{k}}$ belongs to a positive class. And let $\hat{p}(\boldsymbol{x}_{g_{k}})$ be the shorthand writing for $\hat{p}(\tilde{y}=1;\boldsymbol{x}_{g_{k}},f_{k})$.
Condition 1 (Ideal Predicted Probability).
@@
When the predicted probability $\hat{p}(\boldsymbol{x}_{g_{k}})$ is ideal, we have $\hat{p}(\boldsymbol{x}_{g_{k}})=\hat{p}(\tilde{y}=1;\boldsymbol{x}_{g_{k}},f_{k})=\hat{p}(\tilde{y}=1;\boldsymbol{x}_{g_{k}}\in\boldsymbol{D}_{y^{*}=r,g=g_{k}},f_{k})=p(\tilde{y}=1;\boldsymbol{x}_{g_{k}}\in\boldsymbol{D}_{y^{*}=r,g=g_{k}},f_{k})=p(\tilde{y}=1|y^{*}=r,g=g_{k})=p(\boldsymbol{x}_{g_{k}})$. Essentially, ideal predicted probability $\hat{p}(\boldsymbol{x}_{g_{k}})$ match the noise rates $p(\tilde{y}=1|y^{*}=r,g=g_{k})$ corresponding to true label $r$ for $\boldsymbol{x}_{g_{k}}$.
Theorem 1.
@@
For a dataset $\boldsymbol{D}\coloneqq(\boldsymbol{x},\tilde{y})^{n}$ with group- and class-conditional label noise $\pi_{0,{g_{k}}}=p(\tilde{y}=0|y^{*}=1,g=g_{k})<0.5$ and $\pi_{1,{g_{k}}}=p(\tilde{y}=1|y^{*}=0,g=g_{k})<0.5$, if the classifiers $f_{k}$ produce ideal predicted probabilities, then the set of detected mislabeled instances $\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g}$ is a consistent estimator for $\boldsymbol{D}_{\tilde{y}\neq y^{*}|g}$.
@@
While Condition 1 provides a useful sanity check, we now relax this assumption to capture a more nuanced and realistic scenario where the noisy predicted probability $\hat{p}(\boldsymbol{x}_{g_{k}})$ can vary at the instance level. This assumption is well-established and widely adopted in prior studies (e.g. jiang2020beyond ; northcutt2021confident ). We extend this perspective to explicitly incorporate both group- and class-conditional noise structures. Under this extended framework, we show that DeCoLe successfully identifies mislabeled instances in a dataset with label bias and that $\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g}$ remains a consistent estimator for $\boldsymbol{D}_{\tilde{y}\neq y^{*}|g}$.
Condition 2 (Per-instance diffracted (Noisy) Predicted Probability).
@@
The predicted probability $\hat{p}(\boldsymbol{x}_{g_{k}})$ provided by model $f_{k}$ is per-instance diffracted if it follows the relationship $\hat{p}(\boldsymbol{x}_{g_{k}})=p(\boldsymbol{x}_{g_{k}})+\epsilon_{\boldsymbol{x}_{g_{k}}}$ where the noise term $\epsilon_{\boldsymbol{x}_{g_{k}}}$ is drawn from the following distribution: $\epsilon_{\boldsymbol{x}_{g_{k}}}\sim\mathcal{U}[\epsilon_{k}+\text{LB}^{*}_{g_{k}}-p(\boldsymbol{x}_{g_{k}}),\epsilon_{k}-\text{LB}^{*}_{g_{k}}+p(\boldsymbol{x}_{g_{k}})]\hskip 1.99997pt$ when $p(\boldsymbol{x}_{g_{k}})>\frac{1}{2}$; and $\epsilon_{\boldsymbol{x}_{g_{k}}}\sim\mathcal{U}[\epsilon_{k}-\text{UB}^{*}_{g_{k}}+p(\boldsymbol{x}_{g_{k}}),\epsilon_{k}+\text{UB}^{*}_{g_{k}}-p(\boldsymbol{x}_{g_{k}})]$ when $p(\boldsymbol{x}_{g_{k}})<\frac{1}{2}$. Here, $\mathcal{U}$ denotes a uniform distribution, and $\epsilon_{k}=\mathbb{E}_{\boldsymbol{x_{g_{k}}}}[\epsilon_{\boldsymbol{x}_{g_{k}}}]$, where $\epsilon_{\boldsymbol{x}_{g_{k}}}$ represents the deviation from the ideal predicted probability. $\text{LB}^{*}_{g_{k}}$ and $\text{UB}^{*}_{g_{k}}$ denote the value of $\text{LB}_{g_{k}}$ and $\text{UB}_{g_{k}}$ under condition 1.
Theorem 2.
@@
For a dataset $\boldsymbol{D}\coloneqq(\boldsymbol{x},\tilde{y})^{n}$, with group- and class-conditional label noise $\pi_{0,{g_{k}}}=p(\tilde{y}=0|y^{*}=1,g=g_{k})<0.5$ and $\pi_{1,{g_{k}}}=p(\tilde{y}=1|y^{*}=0,g=g_{k})<0.5$, if the classifiers $f_{k}:\boldsymbol{x}_{g_{k}}\rightarrow\hat{p}(\boldsymbol{x}_{g_{k}})$ yields per-instance diffracted predicted probabilities, then the set of detected mislabeled instances $\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g}$ is a consistent estimator for $\boldsymbol{D}_{\tilde{y}\neq y^{*}|g}$.
@@
The theoretical results establish that DeCoLe reliably performs bias-aware mislabeling detection under well-defined conditions. First, when the model produces ideal predicted probabilities, DeCoLe perfectly captures all erroneous labels. Then, relaxing this assumption, we consider the more realistic case where the predicted probabilities are per-instance diffracted (noisy), demonstrating that DeCoLe consistently identifies mislabeled instances. The detailed proofs of these theorems are provided in Appendix B.
@@ 4.1   Evaluation Metrics
As illustrated in Figure 1, DeCoLe partitions the dataset into two disjoint subsets: (1) estimated mislabeled data, denoted by $\boldsymbol{\hat{D}}{\tilde{y}\neq y^{*}|g}$ and (2) estimated correctly labeled data, denoted by $\boldsymbol{\hat{D}}{\tilde{y}=y^{*}|g}$. This decomposition forms the foundation for evaluating DeCoLe’s performance. To quantify performance, we define four evaluation metrics:
@@
First, for data estimated to be mislabeled, we assess the proportion of truly mislabeled instances (denoted by $\boldsymbol{D}_{\tilde{y}\neq y^{*}|g}$) that are successfully identified for scrutiny. This corresponds to the overall recall of mislabeled instances, and is defined in Equation 5.
$\textit{Recall}_{\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g}}=\frac{|\boldsymbol{D}_{\tilde{y}\neq y^{*}|g}\cap\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g}|}{|\boldsymbol{D}_{\tilde{y}\neq y^{*}|g}|}$ (5)
@@
$\textit{Precision}_{\boldsymbol{\hat{D}}_{\tilde{y}=y^{*}|g}}=\frac{|\boldsymbol{D}_{\tilde{y}=y^{*}|g}\cap\boldsymbol{\hat{D}}_{\tilde{y}=y^{*}|g}|}{|\boldsymbol{\hat{D}}_{\tilde{y}=y^{*}|g}|}$ (6)
@@
$\textit{Recall}_{\boldsymbol{\hat{D}}_{\tilde{y}=c,y^{*}=\overline{c}|g}}=\frac{|\boldsymbol{D}_{\tilde{y}=c,y^{*}=\overline{c}|g}\cap\boldsymbol{\hat{D}}_{\tilde{y}=c,y^{*}=\overline{c}|g}|}{|\boldsymbol{D}_{\tilde{y}=c,y^{*}=\overline{c}|g}|}$ (7)
@@
$\textit{Precision}_{\boldsymbol{\hat{D}}_{\tilde{y}=y^{*}=c|g}}=\frac{|\boldsymbol{D}_{\tilde{y}=y^{*}=c|g}\cap\boldsymbol{\hat{D}}_{\tilde{y}=y^{*}=c|g}|}{|\boldsymbol{D}_{\tilde{y}=y^{*}=c|g}|}$ (8)
@@ 4.2.1   Data Generation
We create a synthetic dataset with group and class conditional noise rates, enabling full control over the relationship between observed labels $\tilde{y}$ and latent gold standard labels $y^{*}$. The synthetic population consists of $N=10{,}000$ instances, each associated with covariates $\boldsymbol{X}\in\mathbb{R}^{2}$, a binary group membership $g\in\{0,1\}$, a latent good standard label $y^{*}\in\{0,1\}$, and an observed label $\tilde{y}\in\{0,1\}$. We consider group imbalance, a frequent occurrence in real-world data (mitchell2018prediction, ), by assigning 70% of the population to a predominant group ($g=1$). Additionally, we account for differential subgroup validity (hunter1979differential, ; de2022algorithmic, ), a common phenomenon in which the relationship between covariates and target labels varies across groups. To do this, we draw covariates for different group and class combinations from bi-dimensional normal distributions with different means. We use the same standard deviation for all normal distributions. Further details about the sampling of $\boldsymbol{X}\in\mathbb{R}^{2}$ and $y^{*}\in\{0,1\}$ are provided in Appendix A.
@@
To elucidate the phenomenon of differential subgroup validity and the rationale behind the decoupling component in DeCoLe, we visualize the linear classifiers that would be learned if the data is pooled together vs. if separate linear classifiers are learned for each group. It is easy to see that the classifiers for group $g_{0}$ only (Figure 3 (b)) and for group $g_{1}$ only (Figure 3 (c)) exhibit fundamental dissimilarity, being nearly orthogonal. Therefore, when we fit one classifier for both groups, as depicted in Figure 3 (a), it demonstrates the risks of differential subgroup validity. This phenomenon is closely related to Simpson’s Paradox (simpson1951interpretation, ). Note that the classification accuracy for the linear separator in Figure 3 (a) is notably higher for the majority group compared to the minority group. Furthermore, the classifier in Figure 3 (a) tends to misclassify positive instances in the minority group $g_{0}$ as negative and misclassify negative instances in the majority group $g_{1}$ as positive, reflecting how the classifier may learn and amplify bias in the data labels. If such a classifier is used as a building block of a mislabeling detection algorithm, then such bias will be encoded in it. As we will show in the following subsections, this results in poor mislabeling detection performance when the group-specific nature of predictive relationships and mislabeling patterns is not accounted for. The results also show how this issue is succesfully address by DeCoLe.